# Counting points on character varieties

### v1

@@ 4.6. Palindromic property
Our goal is to explain that the polynomial $|{X}^{F}|$ is palindromic. Recall (cf. [DM20, §7.2]) that the Alvis–Curtis duality is (up to a sign) an involution
@@
To compute the dimension of the irreducible character $\mathfrak{D}(\chi)$, we need to replace $q$ by $q^{-1}$ in the polynomial $d_{\chi}$ encoding the dimension of $\chi$. More precisely,
$\dim\mathfrak{D}(\chi)=q^{|\Phi^{+}|}d_{\chi}(q^{-1}).$
4.6.1.
It is known that $\mathfrak{D}$ restricts to an involution on the set of irreducible constituents of $R_{T}^{G}\,\theta$; namely, it sends the irreducible constituent corresponding to $\rho\in\mathrm{Irr}(W_{\theta})$ to the constituent corresponding to $\epsilon\rho$, where $\epsilon$ is the sign character of $W_{\theta}$. In particular, we see that $\mathfrak{D}$ sends an irreducible character of type $[L,\rho]$ to one of type $[L,\epsilon\rho]$. Using these facts, it is easy to show that $|{X}^{F}|$ is palindromic. For details, see [Whitbread24, §7.7].
4.7. Multi-punctured case
Let ${X}$ be a character variety associated to an $n$-punctured surface group satisfying Assumption 1. Then a similar analysis as in the once-punctured case gives
$|{X}^{F}|=\frac{|Z^{F}|}{|T^{F}|^{n}}\sum_{\tau\in\mathcal{T}(G)}S_{\tau}(q)\,m_{\tau}(q)^{2g-2+n},$
where
$S_{\tau}(q):=\sum_{\chi\in\mathrm{Irr}(G^{F})_{\tau}}\prod_{i=1}^{n}\chi(S_{i})=\frac{\dim(\rho)^{n}\,|W|^{n-1}\,|[L]|\,\nu(L)}{|W(L)|^{n-1}}|Z^{F}|.$
For details, see [Whitbread24, §6.2].
4.7.1.
For the benefit of the reader, we remind what the notation means. Here, $Z$ is the centre and $T$ is a split maximal torus of $G$, $\mathcal{T}(G)$ is the set of $G$-types consisting of $W$-orbits of pairs $(L,\rho)$, where $L$ is a standard endoscopy group for $G$ and $\rho$ is an irreducible character of $W(L)$. The integer $|[L]|$ is the size of the $W$-orbit of $L$, and the integer $\nu(L)$ is defined in §4.4 using the Möbius function on the poset of endoscopy subsystems of $\Phi$. Finally,
$m_{\tau}(q):=q^{|\Phi^{+}(G)|-|\Phi^{+}(L)|}\frac{|L^{F}|}{\dim(\tilde{\rho})}\in\mathbb{Z}[q],$
where $\tilde{\rho}$ is the unipotent character of $L^{F}$ corresponding to $\rho$.
@@ 5. Counting points on additive character varieties
In this section, we count points on the additive character variety introduced in §1.3. We assume first that we have a single puncture and mention the modifications necessary for the multi-punctured case at the end of this section.
5.0.1.
We continue using the notation of the previous section and let $\mathfrak{g}=\mathrm{Lie}({G})$ and $\mathfrak{t}=\mathrm{Lie}(T)$. Let $H\in\mathfrak{t}^{F}$ be a generic regular element and $A$ (resp. $Y$) the additive representation variety (resp. additive character variety) associated to the orbit of $H$ and the once-punctured surface group. By the discussions of §3.4, we have
(6) $|{Y}^{F}|=\frac{|Z^{F}||\mathfrak{g}^{F}|^{g-1}}{|G|^{F}}\sum_{x\in\mathfrak{g}^{F}}|\mathfrak{g}_{x}^{F}|^{g}\mathcal{F}(1_{H}^{G})(x),$
@@ 5.1.1. Green function
5.1.2.
It is known that, if $q$ is large enough, $Q_{T}^{G}(u)$ is a polynomial in $q$ with non-negative coefficients [Springer84, Theorem 5.2]. In fact, as shown in [GoodwinRohrle, Theorem 3.10], under the assumption that $Z$ is connected and $p$ is good for $G$, $Q_{T}^{G}(u)$ is a polynomial in $q$, unless $G$ has a factor of type $E_{8}$, in which case a quadratic base change may be necessary.
Convention 26.
We assume henceforth that $q$ is large enough so that $Q_{T}^{L}(\mathcal{O})$ is a polynomial in $q$ with non-negative coefficients for all standard Levi subgroups $L\subseteq G$ and all unipotent orbits in $\mathfrak{l}$.
@@ 5.1.3. Springer isomorphism
5.1.4.
Let $\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}$ denote the space of $G^{F}$-invariant functions on $\mathfrak{g}^{F}$. For each $x\in\mathfrak{g}^{F}$, let $1_{x}^{G}$ denote the characteristic function of the adjoint orbit $G^{F}.x\subseteq\mathfrak{g}^{F}$. The set ${\{1_{x}^{G}\,|\,x\in\mathfrak{g}^{F}/G^{F}\}}$ is a basis of $\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}$. One can think of this set as the additive analogue of the set of irreducible complex characters of $G^{F}$. The Fourier transform $\mathcal{F}:\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}\rightarrow\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}$ is defined by
@@
5.1.5.
We write $x=x_{s}+x_{n}$ for the Jordan decomposition of $x\in\mathfrak{g}^{F}$. Define the Harish-Chandra induction map
@@
by
$R_{\mathfrak{t}}^{\mathfrak{g}}(f)(x)=\frac{1}{|G_{x_{s}}^{F}|}\sum_{\{g\in G^{F}\,|\,g.x_{s}\in\mathfrak{t}^{F}\}}Q_{T}^{G_{x_{s}}}(\varpi(x_{n}))f(g.x_{s}).$
@@ 5.2. A theorem of Kazhdan and Letellier
5.2.2.
Let $x\in(\mathfrak{g}^{F})^{\circ}$ be a split element. Then $x_{s}$ is $G^{F}$-conjugate to an element $t\in\mathfrak{t}^{F}$, well-defined up to $W$-conjugation. So, the centraliser $G_{x_{s}}$ is a standard Levi subgroup of $G$, well-defined up to the action of $W$. Note that since $x_{n}$ and $x_{s}$ commute, we have $x_{n}\in\mathfrak{g}_{x_{s}}$; thus, we can consider the nilpotent orbit $G_{x_{s}}^{F}.x_{n}\subseteq\mathfrak{g}_{x_{s}}^{F}$.
5.2.3. Types
@@
First of all, since $p$ is assumed to be good for $\mathfrak{g}$, it is also good for $\mathfrak{l}$. In this case, it is known that nilpotent orbits in $\mathfrak{l}$ are in bijection with nilpotent orbits in $\mathfrak{l}(\mathbb{C})$, cf. [Premet03]. Next, for any element $x\in\mathfrak{l}$, consider the orbit $L.x$. The Frobenius fixed point $(L.x)^{F}$ splits into $N_{x}$ many $L^{F}$-orbits, where $N_{x}:=|H^{1}(F,\pi_{0}(L_{x}))|$, cf. [DM20, Proposition 4.2.14]. Finally, observe that $N_{x}$ is the number of $F$-conjugacy classes in $\pi_{0}(L_{x})$. By Corollary 8, $p\nmid|\pi_{0}(L_{x})|$. Thus, the number of $F$-conjugacy classes in $\pi_{0}(L_{x})$ is independent of $k$. ∎
@@
5.2.5.
Let $\tau=[L,\mathcal{N}]$ be a type. Recall that this means that $\mathcal{N}$ is the $L^{F}$-orbit of a nilpotent element $n\in\mathfrak{l}^{F}$. Choose $n\in\mathcal{N}$ and let $d(\mathcal{N}):=\dim(L.n)$. Clearly, this is independent of the choice of $n\in\mathcal{N}$. Let
$d(\tau):=\dim(L)-d(\mathcal{N}).$
The following is the additive analogue of Proposition 20:
@@ 5.3. Evaluating the sum $H_{\tau}$
We first show that the size of the orbit $|\mathfrak{O}|$ depends only on the type of $\mathfrak{O}$:
@@
Let $x=x_{s}+x_{n}\in\mathfrak{g}^{F}$. Jordan decomposition is preserved under adjoint action; thus, for $g\in G$, $g.x=g.x_{s}+g.x_{n}$ is the Jordan decomposition of $g.x$. It follows that $g.x=x$ if and only if $g.x_{s}=x_{s}$ and $g.x_{n}=x_{n}$. In other words,
@@
$\mathcal{A}(x):=\{g\in G^{F}\,|\,g.x_{s}\in\mathfrak{t}^{F}\}.$
If we assume $x_{s}\in\mathfrak{t}^{F}$, then one readily checks
@@
Therefore, under this assumption we have:
$R_{\mathfrak{t}}^{\mathfrak{g}}(f_{H})(x)=\frac{Q_{T}^{G_{x_{s}}}(\varpi(x_{n}))}{|G_{x_{s}}^{F}|}\sum_{g\in\mathcal{A}(x)}f_{H}(g.x_{s})=\frac{Q_{T}^{G_{x_{s}}}(\varpi(x_{n}))}{|W_{x_{s}}|}\sum_{w\in W}f_{H}(w.x_{s}).$
5.3.2.
@@
$H_{\tau}=\frac{|G^{F}||\mathcal{N}|}{|L^{F}|}\sum_{x\in\mathfrak{g}^{F}_{\tau}/G^{F}}{q^{|\Phi^{+}|}}R_{\mathfrak{t}}^{\mathfrak{g}}(f_{H})(x)=\frac{|G^{F}||\mathcal{N}|}{|L^{F}|}\frac{q^{|\Phi^{+}|}Q_{T}^{L}(\varpi(\mathcal{N}))}{|W(L)|}\sum_{w\in W}\,\,\sum_{x\in\mathfrak{g}^{F}_{\tau}/G^{F}}f_{H}(w.x_{s}).$
@@
5.3.3.
The element $x_{s}$ is well-defined up to $W/W_{x_{s}}$-conjugacy, thus, we can now rewrite the above sum as:
$\sum_{w\in W}\sum_{x\in\mathfrak{g}^{F}_{\tau}/G^{F}}f_{H}(w.x_{s})=\frac{|W(L)|}{|W|}\sum_{w\in W}\sum_{{\begin{subarray}{c}y\in\mathfrak{t}^{F}\\ [W_{y}]=[W(L)]\end{subarray}}}f_{H}(w.y)=\frac{|W(L)|}{|N_{W}(W(L))|}\sum_{w\in W}\sum_{{\begin{subarray}{c}y\in\mathfrak{t}^{F}\\ W_{y}=W(L)\end{subarray}}}f_{H}(w.y).$
5.3.4.
@@
$\boxed{H_{\tau}(q)=\frac{q^{|\Phi^{+}|}\,Q_{T}^{L}(\varpi(\mathcal{N}))\,|\mathcal{N}|\,|G^{F}|\,|[L]|}{|L^{F}|\,|W|}\sum_{w\in W}\beta_{L,w.H}(q).}$
Thus, the problem of computing $H_{\tau}$ reduces to evaluating $\beta_{L,H}$. Note that this is the additive analogue of $\alpha_{L,S}$ considered in the previous section.

### v2

@@ 5.1. Palindromic property
We start by recalling some facts about Alvis–Curtis duality; see [DM20, §7.2] for further details.
5.1.1.
Alvis–Curtis duality is (up to a sign) an involution
@@
To compute the dimension of the irreducible character $\mathfrak{D}(\chi)$, we need to replace $t$ by $t^{-1}$ in the polynomial $d_{\chi}(t)\in\mathbb{Q}[t]$ encoding the dimension of $\chi$. More precisely,
$\dim\mathfrak{D}(\chi)=t^{|\Phi^{+}|}d_{\chi}(t^{-1}).$
5.1.2.
It is known that $\mathfrak{D}$ restricts to an involution on the set of irreducible constituents of $R_{T}^{G}\,\theta$; namely, it sends the irreducible constituent corresponding to $\rho\in\mathrm{Irr}(W_{\theta})$ to the constituent corresponding to $\epsilon\rho$, where $\epsilon$ is the sign character of $W_{\theta}$. In particular, we see that $\mathfrak{D}$ sends an irreducible character of type $[L,\rho]$ to one of type $[L,\epsilon\rho]$.
5.1.3.
Now, observe that
$f_{Z}(t^{-1})=(-1)^{\dim(Z)}t^{-\dim(Z)}f_{Z}(t),\qquad f_{T}(t^{-1})=(-1)^{\dim(T)}t^{-\dim(T)}f_{T}(t),$
and
$f_{L}(t^{-1})=(-1)^{\dim(T)}t^{-\dim(T)-3|\Phi^{+}(L)|}f_{L}(t).$
Moreover, we have
$m_{\tau}(t^{-1})=(-1)^{\dim(T)}t^{-\dim(T)-|\Phi(G)|}m_{\tau^{\prime}}(t),$
where $\tau^{\prime}=[L,\epsilon\rho]$. Thus, we conclude:
$f_{X}(t^{-1})=t^{-\deg(f_{X})}\frac{|W|^{n-1}f_{Z}(t)^{2}}{f_{T}(t)^{n}}\sum_{\tau=[L,\rho]\in\mathcal{T}(G)}\frac{\dim(\rho)^{n}|[L]|\,m_{\tau^{\prime}}(t)^{2g-2+n}}{|W(L)|^{n-1}}.$
Since the map $\tau=[L,\rho]\mapsto\tau^{\prime}=[L,\epsilon\rho]$ is an involution, we conclude $t^{\deg(f_{X})}f_{X}(t^{-1})$ equals $f_{X}(t)$; i.e., $f_{X}$ is a palindrome. ∎
5.2. Number of connected components
The approach here is very similar to [BK22, Corollary 5], where it is shown that the number of components of the character stack associated to a compact surface group is $|\pi_{0}(Z(\check{G}))|$. Thus, we only sketch the main idea.
Since $X$ is smooth and equidimensional, connected components and irreducible components coincide. Thus, the number of connected components equals the leading coefficient of the $E$-polynomial. By the above discussions, the $E$-polynomial of $\overline{X}$ equals $f_{X}$. Thus, our goal is to determine the leading term of $f_{X}$.
Following the same analysis as [BK22, Corollary 5], one can show that the only type contributing to the leading term of $f_{X}$ is the $[G,1]$. This is the type corresponding to one-dimensional (principal series) representations of $G(\mathbb{F}_{q^{r}})$; i.e. representations of the form
$G(\mathbb{F}_{q^{r}})\rightarrow G(\mathbb{F}_{q^{r}})^{\mathrm{ab}}\rightarrow\mathbb{C}^{\times}.$
The number of such representations is
$|G(\mathbb{F}_{q^{r}})^{\mathrm{ab}}|=|Z(\check{G}(\mathbb{F}_{q^{r}}))|=|\pi_{0}(\check{Z}(G))(\mathbb{F}_{q^{r}})|\times q^{r\dim(Z)}.$
As $q$ is coprime to $|\pi_{0}(\check{Z}(G))|$ and $r$ is chosen so that $|\pi_{0}(\check{Z}(G))(\mathbb{F}_{q^{r}})|=|\pi_{0}(\check{Z}(G))|$, we conclude that the leading coefficient of $f_{X}$ is
$|\pi_{0}(Z(\check{G}))(\mathbb{F}_{q^{r}})|=|\pi_{0}(Z(\check{G}))|.$
5.3. Euler characteristic
Note that $E(\overline{X})=f_{X}(1)$. We have three cases:
5.3.1. Genus $g=0$
The formula 6 implies
$f_{X}(1)=\frac{1}{(2r)!}\frac{d^{2r}}{dq^{2r}}\bigg{|}_{q=1}\xi(q),$
where $r:=\dim(T)-\dim(Z)$ is the semisimple rank of $G$ and
$\xi(q):=q^{|\Phi^{+}|(n-2)}\sum_{L}\bigg{(}\frac{|W|}{|W(L)|}\bigg{)}^{n-1}\nu(L)\sum_{\rho}\dim(\rho)^{n}\bigg{(}\frac{P_{W(L)}(q)}{\dim(\tilde{\rho})}\bigg{)}^{n-2}.$
Here, the first sum is over all standard endoscopy groups of $G$, and the second sum is over all irreducible characters of $W(L)$.
5.3.2. Genus $g=1$ and trivial centre
The formula (6) implies
$f_{X}(1)=|W|^{n-1}\sum_{L}|W(L)|\,|\mathrm{Irr}(W(L))|\,\nu(L),$
where the sum is over all standard endoscopy groups of $G$.
5.3.3. Genus $g>1$ or $g=1$ and non-trivial centre
We shall prove that $f_{X}(1)=0$, thus establishing Corollary 3 (c). By Remark 26, the polynomial $f_{T}(t)$ divides $m_{\tau}(t)$. This implies that if $g>1$, then $f_{T}$ divides $f_{X}$; hence $f_{X}(1)=0$. It also implies that if $g=1$, then $f_{Z}$ divides $f_{X}$. Since $Z$ is nontrivial and connected, it follows that $f_{Z}(1)=0$. Thus, $f_{X}(1)=0$.
@@ 6. Counting points on additive character varieties
In this section, we count points on the additive character variety introduced in §1.3 and prove Theorem 6 and Corollary 8. For the ease of notation, we assume first that we have a single puncture and mention the modifications necessary for the multi-punctured case at the end of the section.
6.0.1.
We continue using the notation of the previous section and let $\mathfrak{g}:=\mathrm{Lie}({G})$ and $\mathfrak{t}:=\mathrm{Lie}(T)$. Let $H\in\mathfrak{t}^{F}$ be a generic regular element and $A$ (resp. $Y$) the additive representation variety (resp. additive character variety) associated to the $G$-orbit of $H$ and the once-punctured surface group. By the additive analogue of the Frobenius Mass Formula (Theorem 20), we have
(7) $|{Y}^{F}|=\frac{|Z^{F}||\mathfrak{g}^{F}|^{g-1}}{|G^{F}|}\sum_{x\in\mathfrak{g}^{F}}|\mathfrak{g}_{x}^{F}|^{g}\mathcal{F}(1_{H}^{G})(x),$
@@ 6.1.2. Springer isomorphism
6.1.3.
Let $\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}$ denote the space of $G^{F}$-invariant functions on $\mathfrak{g}^{F}$. For each $x\in\mathfrak{g}^{F}$, let $1_{x}^{G}$ denote the characteristic function of the adjoint orbit $G^{F}\cdot x\subseteq\mathfrak{g}^{F}$. The set
${\{1_{x}^{G}\,|\,x\in\mathfrak{g}^{F}/G^{F}\}}$
is a basis of $\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}$. One can think of this set as the additive analogue of the set of irreducible complex characters of $G^{F}$. The Fourier transform $\mathcal{F}:\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}\rightarrow\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}$ is defined by
@@
6.1.4.
We write $x=x_{s}+x_{n}$ for the Jordan decomposition of $x\in\mathfrak{g}^{F}$. Following [Letellier05], the Harish-Chandra induction map
@@
is defined by
$R_{\mathfrak{t}}^{\mathfrak{g}}(f)(x)=\frac{1}{|G_{x_{s}}^{F}|}\sum_{\{g\in G^{F}\,|\,g\cdot x_{s}\in\mathfrak{t}^{F}\}}Q_{T}^{G_{x_{s}}}(\varpi(x_{n}))f(g\cdot x_{s}).$
@@ 6.2. A theorem of Kazhdan and Letellier
6.2.2.
Now, let $x\in(\mathfrak{g}^{F})^{\circ}$ be a split element. Then $x_{s}$ is $G^{F}$-conjugate to an element $t\in\mathfrak{t}^{F}$, well-defined up to $W$-conjugation. So, the centraliser $G_{x_{s}}$ is a standard Levi subgroup of $G$, well-defined up to the action of $W$. Note that since $x_{n}$ and $x_{s}$ commute, we have $x_{n}\in\mathfrak{g}_{x_{s}}$; thus, we can consider the orbit $G_{x_{s}}^{F}\cdot x_{n}\subseteq\mathfrak{g}_{x_{s}}^{F}$.
6.2.3. $\mathfrak{g}$-types
@@
First of all, since $p$ is assumed to be good for $\mathfrak{g}$, it is also good for $\mathfrak{l}$. In this case, it is known that nilpotent orbits in $\mathfrak{l}$ are in bijection with nilpotent orbits in $\mathfrak{l}(\mathbb{C})$, cf. [Premet03]. Next, for any element $x\in\mathfrak{l}$, consider the orbit $L\cdot x$. The Frobenius fixed point $(L\cdot x)^{F}$ splits into $N_{x}$ many $L^{F}$-orbits, where $N_{x}:=|H^{1}(F,\pi_{0}(L_{x}))|$, cf. [DM20, Proposition 4.2.14]. Finally, observe that $N_{x}$ is the number of $F$-conjugacy classes in $\pi_{0}(L_{x})$. By Corollary 11, $p\nmid|\pi_{0}(L_{x})|$. Thus, the number of $F$-conjugacy classes in $\pi_{0}(L_{x})$ is independent of $k$. This concludes the proof of the proposition. ∎
@@
6.2.5.
Let $\tau=[L,\mathcal{N}]$ be a type. Choose $n\in\mathcal{N}$ and let
$d(\tau):=\dim(C_{L}(n)).$
Clearly, this is independent of the choice of $n\in\mathcal{N}$. The following is the additive analogue of Proposition 25:
@@ 6.3. Evaluating the sum $H_{\tau}$
Next, we show that the size of the orbit $|\mathfrak{O}|$ depends only on the type of $\mathfrak{O}$:
@@
Let $x=x_{s}+x_{n}\in\mathfrak{g}^{F}$. Jordan decomposition is preserved under adjoint action; thus, for $g\in G$, $g\cdot x=g\cdot x_{s}+g\cdot x_{n}$ is the Jordan decomposition of $g\cdot x$. It follows that $g\cdot x=x$ if and only if $g\cdot x_{s}=x_{s}$ and $g\cdot x_{n}=x_{n}$. In other words,
@@
$\mathcal{A}(x):=\{g\in G^{F}\,|\,g\cdot x_{s}\in\mathfrak{t}^{F}\}.$
If we assume $x_{s}\in\mathfrak{t}^{F}$, then
@@
Therefore, under this assumption, we have:
$R_{\mathfrak{t}}^{\mathfrak{g}}(f_{H})(x)=\frac{Q_{T}^{G_{x_{s}}}(\varpi(x_{n}))}{|G_{x_{s}}^{F}|}\sum_{g\in\mathcal{A}(x)}f_{H}(g\cdot x_{s})=\frac{Q_{T}^{G_{x_{s}}}(\varpi(x_{n}))}{|W_{x_{s}}|}\sum_{w\in W}f_{H}(w\cdot x_{s}).$
6.3.2.
@@
$H_{\tau}=\frac{|G^{F}||\mathcal{N}|}{|L^{F}|}\sum_{x\in\mathfrak{g}^{F}_{\tau}/G^{F}}{q^{|\Phi^{+}|}}R_{\mathfrak{t}}^{\mathfrak{g}}(f_{H})(x)=\frac{|G^{F}||\mathcal{N}|}{|L^{F}|}\frac{q^{|\Phi^{+}|}Q_{T}^{L}(\varpi(\mathcal{N}))}{|W(L)|}\sum_{w\in W}\,\,\sum_{x\in\mathfrak{g}^{F}_{\tau}/G^{F}}f_{H}(w\cdot x_{s}).$
@@
6.3.3.
The element $x_{s}$ is well-defined up to $W/W_{x_{s}}$-conjugacy, thus, we can now rewrite the above double sum as:
$\sum_{w\in W}\sum_{x\in\mathfrak{g}^{F}_{\tau}/G^{F}}f_{H}(w\cdot x_{s})=\frac{|W(L)|}{|W|}\sum_{w\in W}\sum_{{\begin{subarray}{c}y\in\mathfrak{t}^{F}\\ [W_{y}]=[W(L)]\end{subarray}}}f_{H}(w\cdot y)=\frac{|W(L)|}{|N_{W}(W(L))|}\sum_{w\in W}\sum_{{\begin{subarray}{c}y\in\mathfrak{t}^{F}\\ W_{y}=W(L)\end{subarray}}}f_{H}(w\cdot y).$
6.3.4.
@@
$\boxed{H_{\tau}(q)=\frac{q^{|\Phi^{+}|}\,Q_{T}^{L}(\varpi(\mathcal{N}))\,|\mathcal{N}|\,|G^{F}|\,|[L]|}{|L^{F}|\,|W|}\sum_{w\in W}\beta_{L,w\cdot H}(q).}$
Thus, the problem of computing $H_{\tau}$ reduces to evaluating $\beta_{L,H}$. Note that $\beta_{L,H}$ is the additive analogue of $\alpha_{L,S}$ considered in the previous section.